OpenStudy (anonymous):

Find an equation for the tangent line to the specified graph at the point given. the graph of y=x^2 at (4, 16). I need detailed explaination because I am just not following. Thanks

OpenStudy (anonymous):

your graph is a parabola that runs through the origin, at the point (4,16) we want to find the equation of the tangent line y=mx+b

OpenStudy (anonymous):

because the line is tangent to our graph the slope of the tangent line and the slope at the point on our graph are equal

OpenStudy (anonymous):

take the derivative of the graph and plug in the value of x to find the slope at that point

OpenStudy (anonymous):

y'=2x evaluated at 4 = 8

OpenStudy (anonymous):

\[m=\frac{y1-y}{x1-x}=\frac{16-y}{4-x}\]

OpenStudy (anonymous):

\[\frac{16-y}{4-x}=8\]

OpenStudy (anonymous):

16-y=8(4-x)

OpenStudy (anonymous):

how is that 8?

OpenStudy (anonymous):

16-32+8x=y and your equation is y=8x-16

OpenStudy (anonymous):

m is equal to slope and we no slope is y'=2(4)=8 so we can set them equal to each other

OpenStudy (anonymous):

i dont under stand the inverse. it goes from y = x^2 to just 2x?

OpenStudy (anonymous):

thats the derivative of the function power rule

OpenStudy (anonymous):

d/dx x^n = nx^n-1

OpenStudy (anonymous):

oh yeah you bring it out front. ok sorry.

OpenStudy (anonymous):

its all good its a tricky question and tough to learn from a message board